OpenStudy (anonymous):

how come x^2y'+2xy can be written as (x^2*y)'?? is there something similar we can do with xln(x)y'+y?

OpenStudy (anonymous):

find the derivative of x^2y using the product rule for differentiation

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

the product rule: \[\frac{ d }{ dx }=f(x)g'(x)+g(x)f'(x)\]

OpenStudy (anonymous):

sorry should be \[\frac{ d }{ dx }(f(x)g(x))=f(x)g'(x)+g(x)f'(x)\]

OpenStudy (anonymous):

so the derivative of x^2y = 2xy+y'x^2 OHHHHHHH

OpenStudy (anonymous):

yeah i said the same thing when it clicked for me LOL

OpenStudy (anonymous):

and no you can't do it for the second one, you'll have to find the integrating factor

OpenStudy (anonymous):

nice! is there a similar relation in xln(x)y'+y. (i am assuming there is because all the questions in that section seem to have these patterns)

OpenStudy (anonymous):

OK!

OpenStudy (anonymous):

well you'll have to find the integrating factor which transforms a linear equation into one that we can solve in that form... can i show you an easier problem than that one to start with, its pretty nasty

OpenStudy (anonymous):

wait, you might want the complete question. the question is xln(x)y'+y=xe^x

OpenStudy (anonymous):

do you see anything i should be doing in this question

OpenStudy (anonymous):

yeah its still nasty lol how about i show you an easy one then get into a nasty one

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

you need to divide by x(ln(x))

OpenStudy (anonymous):

i can show you this one but its gonna be ugly

OpenStudy (anonymous):

When we combine the Chain and Product Rules, we get:\[\bf (x^2y)'=(x^2)'y+x^2(y)'=2xy+x^2\frac{ dy }{ dx }\]

OpenStudy (anonymous):

@natto

OpenStudy (anonymous):

yup i get y'+y/xlnx = xe^x/xlnx

OpenStudy (anonymous):

i treid isolating y from x, so i can assign dy and dx to them respectively, but clearly i wasn't able to

OpenStudy (anonymous):

ok its in linear form like this

OpenStudy (anonymous):

its gonna take a bit for me to type

OpenStudy (anonymous):

alright! take ur time!

OpenStudy (anonymous):

\[y'+\frac{ 1 }{ xln(x) }y=\frac{ e^{x} }{ \ln(x) }\]

OpenStudy (anonymous):

ok what we need is a magic function to multiply both sides of the equation so we can use that fancy product rule. Turns out there is a trick to find this function.

OpenStudy (anonymous):

oh nm that integral isn't that bad ha! it just looks bad

OpenStudy (anonymous):

ok if the equation is of the form: \[y'+P(x)y=f(x)\] then the magic function is found by \[e ^{\int\limits_{}^{}P(x)dx}\]

OpenStudy (anonymous):

can you find \[\int\limits_{}^{}\frac{ 1 }{ xln(x) }dx\] (hint: its a u substitution)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok get that real quick

OpenStudy (anonymous):

ln(ln(x))+c

OpenStudy (anonymous):

good ok now what is \[e ^{\ln(\ln(x))}\]

OpenStudy (anonymous):

remember inverses (don't worry about abs bars for now, this isn't nitpick class)

OpenStudy (anonymous):

e^1

OpenStudy (anonymous):

its just ln(x)

OpenStudy (anonymous):

remember e^ln(whatever) = whatever

OpenStudy (anonymous):

ok!!

OpenStudy (anonymous):

Now that we found our magical function, we're gonna multiply both sides of the equation by it, and you will see it become much prettier (i've done this problem before, we must have the same book)

OpenStudy (anonymous):

\[(\ln(x))y'+\frac{ 1 }{ x }y=e^{x}\]

OpenStudy (anonymous):

the left side is now the product rule... poof majic

OpenStudy (anonymous):

see? the left hand side is the derivative of (ln(x))y

OpenStudy (anonymous):

omg how do you see that far ahead ?

OpenStudy (anonymous):

i've taught differential equations 6 semesters :) you see the pattern one is the derivative of the other

OpenStudy (anonymous):

that and its magic... it always works.. these are always solveable

OpenStudy (anonymous):

if you multply a derivative in standard linear form by that magic e^integral of P it always gives you a product rule... always

OpenStudy (anonymous):

but we're not done we need to solve for y...

OpenStudy (anonymous):

Oh! is there a name for that? or more examples that i can up on? i will look it up!

OpenStudy (anonymous):

now we see the product rule on the left we have \[\frac{ d }{ dx }[\ln(x)y]=e^{x}\]

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

you said "magic e^integral of P it always gives you a product rule... always" is there a name or keyword for that rule so i can search it up and read on it later?

OpenStudy (anonymous):

you need to antidifferentiate both sides using the Fundamental Theorem of Calculus... yes i give you information when we finish the problem :)

OpenStudy (anonymous):

ok, working ahead , i get y=e^x/(ln(x))

OpenStudy (anonymous):

almost yu forgot the constant

OpenStudy (anonymous):

right! +c!! e^x+c/ln(x)

OpenStudy (anonymous):

\[y=\frac{ e^{x} }{ \ln(x) }+\frac{C}{\ln(x)}\]

OpenStudy (anonymous):

good job... These are called Linear First order Ordinary Differential Equations. They for the most part are solved using the same method. The "majic equation" is called an integrating factor. This is a good video by a professor at MIT here http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-3-solving-first-order-linear-odes/

OpenStudy (anonymous):

thank you so much!! you've been super helpful!

OpenStudy (anonymous):

np. Differential equations are fun especially when student has a good background in calculus like yourself. :)

OpenStudy (anonymous):

ill watch the video and go to sleep :D thanks alot ! good night!